OpenStudy (theeric):

Hi! I'm having trouble with an ODE. I can solve for solutions in the form \(x^2y''+\alpha y' +\beta y=0\) But now I have \((x-1)^2y''+8(x-1)y' +12 y=0\) with the solution \(y=\dfrac{c_1}{(x-1)^3}+\dfrac{c_2}{(x-1)^4}\)

OpenStudy (theeric):

I considered a substitution, but then (I probably used illegal moves, but) \(\dfrac{d^2y}{du^2}\) where \(u=x-1\) is undefined.

OpenStudy (isaiah.feynman):

You're a better physicist than me Eric. Can't help here.

OpenStudy (loser66):

power series method

OpenStudy (theeric):

Haha, thanks. It's hard to compare physics knowledge! Anyway, this is mathematics!

OpenStudy (theeric):

Thank you @Loser66 , I'll look into that!

OpenStudy (anonymous):

what's wrong with u = x - 1?

OpenStudy (theeric):

For series solutions, don't they need to be around a point? Can I just assume the point where \(x=0\) has a radius of convergence of \(\infty\)? The problem indicates that the solution doesn't need to work at the singular points.

OpenStudy (loser66):

nope, regular singular point is x =1 because it makes y"=0

OpenStudy (theeric):

If I use \(u=x-1\), then I can't have \(\dfrac{\text dy}{\text dx}\) and still solve it with the shortcut \(y=c_1x^{r_1}+c_2x^{r_2}\) where \(r_1\) and \(r_2\) are the roots of \(r^2+(\alpha -1) r+\beta=0\) assuming two distinct real roots?

OpenStudy (theeric):

The singular point is \(x=1\), yep. But I just need to find a solution for all \(x\neq1\).

OpenStudy (theeric):

So, I guess I can try the power series solution about \(x=0\) and suppose that it would converge for all values that aren't \(x=1\)?

OpenStudy (theeric):

\(\dfrac{\text dy}{\text du}=\dfrac{\text dy}{\text dx}\dfrac{\text dx}{\text du}\) where \(\dfrac{\text du}{\text dx}=\dfrac{\text d}{\text dx}\left(x-1\right)=1\)\(\implies\dfrac{\text dx}{\text du}=1\) I don't know if that's allowed... Either way, \(u+1=x\), so \(\dfrac{\text dx}{\text du}=1\) So then \(\dfrac{\text dy}{\text du}=\dfrac{\text dy}{\text dx}\dot\ 1=y'\) That works. But a second derivative would cause \(\dfrac{\text d^2y}{\text du}=0\), I think. I erased what I had. I'll seek a power series solution, I suppose.

OpenStudy (theeric):

Haha, just now found \((x-1)^2y''\). The next ones should be quicker.

OpenStudy (loser66):

i think you can handle from this

OpenStudy (theeric):

Thank you :)

OpenStudy (loser66):

the rest is easy, right? hehehe

OpenStudy (theeric):

Yeah! Thank you very much! I'll get to that point on my own and continue from there. When I take the derivative of the series, though, I shift the index to eliminate the term that became zero upon differentiating.

OpenStudy (theeric):

And then I shift indeces so that all \(x\) are to the \(n\) power, but, now that I think about it, that just makes more work.

OpenStudy (theeric):

Or maybe it's necessary. I can't be sure just by looking at it. But I have to much work to do for me to use my time understanding what will and won't work, unfortunately.

OpenStudy (loser66):

My prof said that if you want to make them have the same form, just convert all of the indices in a sum, for example in y", its sum is \[\sum_{n=0}^{\infty} n(n-1) a_n x^{n-2}\] if you want all indices = n, then, replace them ALL by +2 into the indices \[\sum_{n=0}^{\infty} (n+2)(n+1) a_{n+2} x^{n}\] Dat sit

OpenStudy (loser66):

so that you can factor x^n out, and the leftover is a power series of coefficients

OpenStudy (theeric):

Right. That's what I do. I also do this: Assume that \(y\) is a power series solution about \(x=0\). Then \[y=\sum_{n=0}^\infty a_nx^n\] and \[\dfrac{\text dy}{\text dx}=\sum_{n=1}^\infty a_nnx^{n-1}=\sum_{n=0}^\infty a_{n+1}(n+1)x^{n}\] and \[\dfrac{\text d^2y}{\text dx^2}=\dfrac{\text dy}{\text dx}\left[\sum_{n=1}^\infty a_nnx^{n-1}\right]=\sum_{n=2}^\infty a_nn(n-1)x^{n-2}=\sum_{n=0}^\infty a_{n+2}(n+2)(n+1)x^n\] I change the starting index according to the derivatives I've taken. But I don't think that is necessary.

OpenStudy (theeric):

Rather, I just don't know.

OpenStudy (theeric):

I didn't reach a recursive relation... I got that \(a_0\), \(a_1\), and \(a_2\) are all dependent on each other, and \(a_n=0\) for \(n\ge2\). So \(\forall n\), \(a_n=0\). Bleh...

OpenStudy (theeric):

Thank you for your help, everyone! I don't need to finish this, so maybe I'll try to tackle it later!